OpenStudy (anonymous):

The following 2 problems you are given x' = Ax + f. Give the FORM of a particular solution that you get from the technique of undetermined coefficients. You do not need to solve.

OpenStudy (anonymous):

a) A = [ 0 1 ] f(t)= [ cos(t) ] [ -1 0 ] [ sin(2t) ]

OpenStudy (anonymous):

I am having trouble understanding what the book says about the different forms for the particular solution.

OpenStudy (anonymous):

I think this one would be xp = Acos(t) + Bsin(t) + Csin(2t) + Dcos(2t)

OpenStudy (anonymous):

b) A = [ 0 1 ] f(t)= [ e^(2t) ] [ 1 0 ] [ e^(-t) ]

OpenStudy (anonymous):

I think this one is xp = Ae^(2t) + Be^(-t)

OpenStudy (anonymous):

ex) if f(t) = col(t,e^t,t^2) xp(t) = At^2 + Bt + c + De^t

zepdrix (zepdrix):

Hmm I don't understand this really... Are you finding eigenvalues and all that kinda stuff? D: Orrrrr..?

OpenStudy (anonymous):

yes, the part I'm studying now is about matrices and eigenvalues. This is part of an initial value problem where we just had to setup the particular solution. I left this 20 pt problem blank on my test :/

zepdrix (zepdrix):

Oh my :O

OpenStudy (anonymous):

I'm going back through it trying to solve the ones I got wrong, but I don't have the answer, makes it frustrating

OpenStudy (anonymous):

@zepdrix this is in an ODE course.

OpenStudy (anonymous):

@ChmE I'll have to look back into how to solve this... let me go find my book. :-p

OpenStudy (anonymous):

@Hero @phi @jim_thompson5910 Can you check my answers?

OpenStudy (anonymous):

thx @oldrin.bataku